an economical automobile can travel 442 miles on just 13 gallons of gasoline.how many gallons of gasoline will be needed for a trip of 1530 miles
Answer by stanbon(75887)

an economical automobile can travel 442 miles on just 13 gallons of gasoline.how many gallons of gasoline will be needed for a trip of 1530 miles
-------
x/1530 = 13/442
-----
x = (13/442)*1530
-----
x = (1/34)*1530
--
x = 45 gallons
================================
Cheers,
Stan H.
=========